OpenStudy (anonymous):

write an absolute value epuation that has 2 and 6 as solutions

OpenStudy (anonymous):

Welcome to OpenStudy, @brenna00 ! If you have any questions about the site and how to get around it, don't hesitate to ask. First things first for this question: Do you know what an absolute value is?

OpenStudy (anonymous):

well I know what one looks like but I don't know the definition

OpenStudy (anonymous):

basically it means that instead of looking at it as positive or negative, you look at how far AWAY it is from 0. So the absolute value of -3 is just 3. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So if you apply that to equations, it would be like this: \[\left| x-10 \right| =6\] Solution 1: x-10=6, so x=16 Solution 2: x-10 = -6, so x =4 The two solutions in that case are 4 and 16. In this case, you need to find two equations whose solutions are 2 and 6.

OpenStudy (anonymous):

(basically you take the absolute value expression and set it equal to a negative number and a positive number, and it gives you two different answers)

OpenStudy (anonymous):

That's fair ;) Let me see if i can help

OpenStudy (anonymous):

How about \[\left| x-4 \right| = 2\] ? Would that work?

OpenStudy (anonymous):

yes it works

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem! :)